ADNAN MUHAMMED

Grade 12,

If X = 9 – 4sqrt5, find (A) X+ 1/X (B) sqrtX + 1/sqrtX

If X = 9 – 4sqrt5, find (A) X+ 1/X
                                   (B) sqrtX + 1/sqrtX 

Grade:9

1 Answers

Arun
5 years ago
Dear student
 
1/x = 1/ 9 – 4 sqrt 5
 = 9 + 4 sqt 5 / (81 – 80) = 9 + 4 sqrt 5
 
x + 1/ x = 18
 
please ask only one question in one thread.